# Data-informativity conditions for structured linear systems with implications for dynamic networks

### v2

@@ 4.2 Including a network structure in the generation of input signals
In line with the notation used in dynamic networks (see e.g. Van den Hof et al. (2013)), we will consider the input $u$ to be constructed as a multivariable filtered version of an external $n_{x}$-dimensional vector signal $x$, such that
$u(t)=G_{u}(q)u(t)+R_{u}x(t)$ (15)
with $G_{u}$ a proper and stable transfer function matrices, with $G_{u}$ being hollow, meaning that $[G_{u}]_{kk}=0$ for all $k$, and $R_{u}$ being composed of a (sub)set of columns of the $m\times m$ identity matrix. Additionally and without loss of generality it is assumed that $\Phi_{x}(\omega)$ is diagonal and $\Phi_{x}(\omega)\succ 0$ for all $\omega$.
@@
Consider a network graph $\mathcal{G}$ with vertex set $\mathcal{V}:=\{u_{k}\}_{k=1,\cdots m}\cup\{x_{k}\}_{k=1,\cdots n_{x}}$ and directed edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$:
  • $(u_{k},u_{\ell})\in\mathcal{E}$ if and only if $[G_{u}]_{\ell k}\neq 0$;
  • $(x_{k},u_{\ell})\in\mathcal{E}$ if and only if $[R_{u}]_{\ell k}\neq 0$.
Vertex $u_{\ell}$ is an out-neighbour of $u_{k}$ if $(u_{k},u_{\ell})\in\mathcal{E}$.
An example of a network graph that generates an input vector $u$ is shown in Figure 1, in the particular situation of a target module $G_{j1}$.
Figure 1: Five node example with target module $G_{j1}(q)$, and dependence structure among the input components $u$ as in (15).
@@
A vertex set $\mathcal{D}^{c}$ is called a disconnecting set from $\mathcal{V}_{1}$ to $\mathcal{V}_{2}$, if upon removal of $\mathcal{D}^{c}$ from the graph, there is no directed path from $\mathcal{V}_{1}$ to $\mathcal{V}_{2}$. It is a minimum disconnecting set if it has the minimum cardinality among all such disconnecting sets.
@@
As a next step we are going to decompose the input signals in $\kappa_{u}^{[j]}$ in three different components.
@@
Let $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ be defined as a minimum disconnecting set from $u_{i}$ to the other components in $\kappa_{u}^{[j]}$, under the constraints that $i\notin\mathcal{D}^{c}$.
The interpretation of this disconnecting set is that it disconnects the input $u_{i}$ of the target module from all other inputs to parametrized modules. To illustrate this, consider the situation as depicted in Figure 1. The disconnecting set $\{u_{2}\}$ disconnects $u_{1}$ from $u_{3}$. As a result, the dependency of $u_{3}$ on $u_{1}$ is fully covered by the signal $u_{2}$. This phenomenon is closely related to the concept of blocked parallel paths from $u_{1}$ to $y_{j}$, i.e. parallel to the target module $G_{j1}$ (Dankers et al., 2016). I.e. $u_{2}$ blocks all connections from $u_{1}$ to $y_{j}$ that are not passing through $G_{j1}$.
@@
The full input vector $u$ is then composed of the elements in $\kappa_{u}^{[j]}$ complemented with the input signals $u_{\!{\scriptscriptstyle\mathcal{F}}}$ related to a priori known terms in $G(q,\theta)_{j*}$.
In the sequel the following set of external signals will appear to be instrumental.
@@
Let $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ be defined as those components in vector signal $x$, that in the graph of the input signal $u$, have a path to $u_{\!{\scriptscriptstyle\mathcal{T}}}$ without passing through $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$.
@@
The expression (17) for $u_{\!{\scriptscriptstyle\mathcal{T}}}$ actually defines the set of external signals $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$, as those external signals that have a path to $u_{\!{\scriptscriptstyle\mathcal{T}}}$ without passing through $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$. Note that for external signals that pass through $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ before reaching $u_{\!{\scriptscriptstyle\mathcal{T}}}$, their contribution to $u_{\!{\scriptscriptstyle\mathcal{T}}}$ is covered by the transfer function $T_{s}(q)$ in (17).
@@
  • Let $u_{(i\cup\!{\scriptscriptstyle\mathcal{D}}^{c})\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$ be the projection of signals $u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$ onto the orthogonal complement of $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$.
@@
    It is not the full input vector $\kappa_{u}^{[j]}$ of inputs that have a parametrized map to the output, that requires excitation. Only a subset of input signals, present in $u_{\{i\}\cup\mathcal{D}^{c}}$ need excitation.
@@
    For exciting this subset of inputs, all external signals $x$ can be used, except for the $x$ signals that are present in $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$. For excitation of $u_{\{i\}\cup\mathcal{D}^{c}}$ we can use those signals $x$ that in the graph of the input network reach $u_{\{i\}\cup\mathcal{D}^{c}}$ without passing through $u_{\!{\scriptscriptstyle\mathcal{T}}}$.
@@
In this section, the node set $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ is defined as a minimum disconnecting set from $u_{i}$ to $\kappa_{u}^{[j]}\backslash\{u_{i}\}$. It is obvious, however, that the presented results hold true for any such disconnecting set. The advantage of selecting a minimum disconnecting set, is that it leads to the smallest number of excitation signals.
Interpretation (a.) above confirms the earlier insights of Dankers et al. (2016), that for consistently estimating a target module $G_{ji}$ it would be sufficient to measure input $u_{i}$ and those inputs that block all parallel paths from $u_{i}$ to $y_{j}$ This has been formalized in the Parallel Path and Loop condition (Dankers et al., 2016), or alternatively phrased as a separating/disconnecting set condition (Dankers et al., 2016; Shi et al., 2022).
The current results show that we do not need to exclude the remaining inputs from the predictor model, as was done in Dankers et al. (2016). We can include the inputs $u_{\!{\scriptscriptstyle\mathcal{T}}}$ in our predictor model, but the concerned input signals do not require additional excitation. In this setting the current result improves on the results of Van den Hof and Ramaswamy (2020); Van den Hof et al. (2023) where persistence of excitation of all inputs was chosen in order to guarantee data-informativity.
@@ 5 Example
We illustrate the result of Theorem 4 through the four-input one-output example as indicated in Figure 1. In the considered situation $G_{j1}$ is the target module to be estimated, while the predictor model is given by:
@@
First we consider the situation that all four modules $G_{jk}(q,\theta)$, $k=1,\cdots 4$ are parametrized.
A minimum disconnecting set $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ from input $u_{i}=u_{1}$ to all other inputs of parametrized modules is $u_{2}$. In an alternative interpretation it is clear that $u_{2}$ blocks all parallel paths (i.e. parallel to the target module $G_{j1}$) from $u_{i}$ to output $y_{j}$.
The signal $u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$ is composed of $u_{i}=u_{1}$, $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=u_{2}$, while $u_{\!{\scriptscriptstyle\mathcal{T}}}=(u_{3},u_{4})$. Possible excitation or noise signals $x_{3}$ and $x_{4}$ will serve as $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$. Since $u$ and $e$ are considered to be independent, and $e$ is white noise, the condition (18) can be replaced by a positive definite spectrum of $u_{(1\cup 2)\perp(x_{3},x_{4})}$.
@@
$u_{1}$ $=$ $\sum_{k=1}^{4}F_{1k}(q)x_{k}(t);$ (21)
$u_{2}$ $=$ $\sum_{k=1}^{4}F_{2k}(q)x_{k}(t).$ (22)
A positive definite spectrum of $u_{(1\cup 2)\perp(x_{3},x_{4})}$ is obtained if the matrix
$\begin{bmatrix}F_{11}&F_{12}\\ F_{21}&F_{22}\end{bmatrix}$
has full row rank, and $x_{1}$ and $x_{2}$ are persistently exciting.
@@
Moreover the signals $x_{3}$ and $x_{4}$ cannot be used to excite $(u_{1},u_{2})$ since they pass through $u_{\!{\scriptscriptstyle\mathcal{T}}}$ before reaching $u_{2}(=u_{\!{\scriptscriptstyle\mathcal{D}}^{c}})$, and therefore they belong to $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ To understand why $x_{3}$ and $x_{4}$ cannot be used for excitation of $(u_{1},u_{2})$ we consider the prediction error
$\varepsilon(t,\theta)=H(q,\theta)^{-1}\left[\sum_{k=1}^{4}\Delta G_{jk}(q,\theta)u_{k}(t)+H^{0}(q)e(t)\right]$ (23)
Writing $u_{k}$ as a function of the variables $x$, this can be rewritten as
$\varepsilon(t,\theta)=H(q,\theta)^{-1}\left[\sum_{k=1}^{4}T_{k}(q,\theta)x_{k}(t)+H^{0}(q)e(t)\right]$ (24)
@@
$T_{1}(q,\theta)$ $=$ $\Delta G_{j1}+\Delta G_{j2}\frac{G_{21}}{1-G_{32}G_{23}}\!+\!\Delta G_{j3}\frac{G_{32}G_{21}}{1-G_{32}G_{23}}$
$T_{2}(q,\theta)$ $=$ $\Delta G_{j2}\frac{1}{1-G_{32}G_{23}}+\Delta G_{j3}\frac{G_{32}}{1-G_{32}G_{23}}$
$T_{3}(q,\theta)$ $=$ $\Delta G_{j2}\frac{G_{23}}{1-G_{32}G_{23}}+\Delta G_{j3}\frac{1}{1-G_{32}G_{23}}$
$T_{4}(q,\theta)$ $=$ $\Delta G_{j2}\frac{G_{23}G_{34}}{1-G_{32}G_{23}}\!+\!\Delta G_{j3}\frac{G_{34}}{1-G_{32}G_{23}}\!+\!\Delta G_{j4}$
where $\Delta G_{jk}:=G_{jk}^{0}(q)-G_{jk}(q,\theta)$, for $k=1,\cdots 4$. The minimum of the cost function $\bar{\mathbb{E}}\varepsilon^{T}(t,\theta)\varepsilon(t,\theta)$ is achieved if the $T_{k}$-dependent terms are zero. Presence of $x_{k}$ implies that $T_{k}(q,\theta)=0$ in the minimum of the cost function. If $x_{1}$ and $x_{2}$ are present and persistently exciting, then $T_{1}$ and $T_{2}$ are forced to be zero, which implies that $\Delta G_{j1}=0$, and thus a consistent estimate of $G^{0}_{j_{1}}$ is guaranteed. This follows from the observation that $T_{1}(q,\theta)=\Delta G_{j1}+G_{21}T_{2}(q,\theta)$. If we would replace $x_{2}$ by presence of $x_{3}$, a similar result does not exist. Introducing $x_{4}$ shows that $T_{4}(q,\theta)$ gets involved in the expression and a new unknown term $\Delta G_{j4}$ is introduced also. This explains that forcing $T_{1}$ and $T_{4}$ to be zero is not sufficient for arriving at $\Delta G_{j1}=0$, or in other words: exciting $x_{1}$ and $x_{4}$ does not lead to a consistent estimate of the target module.
In the case that module $G_{j2}(q)$ is not parametrized and known a priori, there are two options for choosing minimum disconnecting sets:
  • When selecting $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=u_{2}$, it follows that $u_{\!{\scriptscriptstyle\mathcal{T}}}=(u_{3},u_{4})$, and as a result $x_{3}$ and $x_{4}$ cannot be used for excitation of $(u_{1},u_{2})$, and therefore we need excitations $x_{1}$ and $x_{2}$.
@@
Note that in case multiple disconnecting sets are possible, each one of them provides a set of sufficient conditions for data-informativity. Not satisfying the conditons for one particular choice of disconnecting set, does not necessarily add to the conclusion that the data-informativity conditions are not satisfied.
In the case that $G_{j4}$ is not parametrized and known a priori, the minimum disconnecting set from $u_{1}$ to $(u_{2},u_{3})$ is given by $u_{2}$. Then $u_{\!{\scriptscriptstyle\mathcal{T}}}=u_{3}$, while the external signals $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ that affect $u_{\!{\scriptscriptstyle\mathcal{T}}}$ are given by $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}=(x_{3},x_{4})$. This implies that (only) signals $(x_{1},x_{2})$ are available for realizing a positive definite spectrum of $(u_{1},u_{2})_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$. Verification of this through the expressions of $T_{k}(q,\theta)$, shows that with $\Delta G_{j4}=0$, the linear combination of the two terms in $T_{3}$ and $T_{4}$ are different from the linear combinations of $\Delta G_{j2}$ and $\Delta G_{j3}$ in the expressions for $T_{1}$ and $T_{2}$. As a result the original result with all modules parametrized follows as the solution.

### v3

@@ 5.2 Including a network structure in the generation of input signals
In line with the notation used in dynamic networks (see e.g. VandenHof&etal_Autom:13 ), we will consider the input $u$ to be constructed as a multivariable filtered version of an external $n_{x}$-dimensional vector signal $x$, such that
$u(t)={\color[rgb]{0,0,0}P}(q)u(t)+{\color[rgb]{0,0,0}Q}(q)x(t)$ (13)
with $P$ a $m\times m$ proper and stable transfer function matrices, with (IP(q))(I-P(q)) invertible, $P$ being hollow, meaning that ${\color[rgb]{0,0,0}P(q)}]_{kk}=0$ for all $k$, and $\color[rgb]{0,0,0}Q$} being composed of a (sub)set of columns of the $m\times m$ identity matrix. Furthermore, and without loss of generality, it is assumed that $\Phi_{x}(\omega)$ is diagonal and $\Phi_{x}(\omega)\succ 0$ for all $\omega$. An example of a network graph that generates an input vector $u$ is shown in Figure 2, in the particular situation of a target module $G_{j1}$.
@@
Consider a network graph $\mathcal{G}$ with vertex set $\mathcal{V}:=\{u_{k}\}_{k=1,\cdots m}\cup\{x_{k}\}_{k=1,\cdots n_{x}}$ and directed edge/link set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$:
  • $(u_{k},u_{\ell})\in\mathcal{E}$ if and only if ${\color[rgb]{0,0,0}P}]_{\ell k}\neq 0$;
  • $(x_{k},u_{\ell})\in\mathcal{E}$ if and only if ${\color[rgb]{0,0,0}Q}]_{\ell k}\neq 0$.
Vertex $u_{\ell}$ is an out-neighbour of $u_{k}$ if $(u_{k},u_{\ell})\in\mathcal{E}$. $\Box$
Figure 2: Five node example with target module $G_{j1}(q)$, and dependence structure among the input components $u$ as in (13), where $v_{j}$ is the $j$-th component of $H^{0}(q)e(t)$ in (1).
@@
A vertex set $\color[rgb]{0,0,0}\mathcal{V}_{\mathcal{D}^{c}}$} is called a disconnecting set from $\mathcal{V}_{1}$ to $\mathcal{V}_{2}$, if upon removal of $\color[rgb]{0,0,0}\mathcal{V}_{\mathcal{D}^{c}}$} from the graph, there is no directed path from $\mathcal{V}_{1}$ to $\mathcal{V}_{2}$. It is a minimum disconnecting set if it has the minimum cardinality among all such disconnecting sets. $\Box$
@@
As a next step we are going to decompose the input signals in $\kappa_{u}^{[j]}$ in three different components. In doing so we adopt the notation that for an $m$-dimensional node vector $u$, and $\mathcal{D}^{c}\subset[1,\cdots m]$ an index set, $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ is a vector of those components of $u$ that are defined by the index set $\mathcal{D}^{c}$. We will use this general notation throughout the remaining part of the paper.
@@
[…]
The interpretation of this disconnecting set is that it disconnects the input $u_{i}$ of the target module from all other inputs to parametrized modules. To illustrate this, consider the situation as depicted in Figure 2. The disconnecting set $\{u_{2}\}$ disconnects ${\color[rgb]{0,0,0}\{}u_{1}{\color[rgb]{0,0,0}\}}$ from ${\color[rgb]{0,0,0}\{}u_{3}{\color[rgb]{0,0,0}\}}$. As a result, the dependency of $u_{3}$ on $u_{1}$ is fully covered by the signal $u_{2}$. This phenomenon is closely related to the concept of blocked parallel paths from $u_{1}$ to $y_{j}$, i.e. parallel to the target module $G_{j1}$ (Dankers&etal_TAC:16, ), i.e. $u_{2}$ blocks all connections from $u_{1}$ to $y_{j}$ that are not passing through $G_{j1}$.
@@
The full input vector $u$ is then composed of the elements in $\kappa_{u}^{[j]}$ complemented with the input signals $u_{\!{\scriptscriptstyle\mathcal{F}}}$ related to a priori known terms in $G(q,\theta)_{j*}$. In the sequel the following set of external signals will appear to be instrumental.
@@
Let $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ be defined as those components in vector signal $x$, that in the graph of the input signal $u$, have a path to $u_{\!{\scriptscriptstyle\mathcal{T}}}$ without passing through $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$. $\Box$
@@
Note that for external signals that pass through $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ before reaching $u_{\!{\scriptscriptstyle\mathcal{T}}}$, their contribution to $u_{\!{\scriptscriptstyle\mathcal{T}}}$ is covered by the transfer function $T_{s}(q)$ in (15).
@@
  • Let $u_{(i\cup\!{\scriptscriptstyle\mathcal{D}}^{c})\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$ be the projection [^2] Projection is considered with respect to the inner product defined by $x,y\!>:=\bar{\mathbb{E}}[x(t)y^{T}(t)]$. of signals $u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$ onto the orthogonal complement of $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$.
@@
    It is not the full input vector $\kappa_{u}^{[j]}$ of inputs that have a parametrized map to the output, that requires excitation. Only a subset of input signals, present in $u_{i\cup\mathcal{D}^{c}}$ need excitation.
@@
    For exciting this subset of inputs, all external signals $x$ can be used, except for the $x$ signals that are present in $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$. For excitation of $u_{i\cup\mathcal{D}^{c}}$ we can use those signals $x$ that in the graph of the input network reach $u_{i\cup\mathcal{D}^{c}}$ without passing through $u_{\!{\scriptscriptstyle\mathcal{T}}}$.
@@
[…]
Interpretation (a.) above confirms the earlier insights of Dankers&etal_TAC:16 , that for consistently estimating a target module $G_{ji}$ it would be sufficient to measure input $u_{i}$ and those inputs that block all parallel paths from $u_{i}$ to $y_{j}$ This has been formalized in the Parallel Path and Loop condition (Dankers&etal_TAC:16, ), or alternatively phrased as a separating/disconnecting set condition (Dankers&etal_TAC:16, ; Shi&etal_Autom:22, ).
The current results show that we do not need to exclude the remaining inputs from the predictor model, as was done in Dankers&etal_TAC:16 . We can include the inputs $u_{\!{\scriptscriptstyle\mathcal{T}}}$ in our predictor model, but the concerned input signals do not require additional excitation. Their effect on output $y_{j}$ is modeled, but for this we do not require a consistent estimate of the concerned transfer functions that map the input signals to $y_{j}$. In this setting the current result improves on the results of VandenHof&Ramaswamy_CDC:20 ; VandenHof&etal_IFAC:23 where persistence of excitation of all inputs was chosen in order to guarantee data-informativity.
@@ 6 Example
We illustrate the result of Theorem 4 through the four-input one-output example as indicated in Figure 2. In the considered situation, $G_{j1}$ is the target module to be estimated, while the predictor model is given by:
@@
First we consider the situation that all four modules $G_{jk}(q,\theta)$, $k=1,\cdots 4$ are parametrized. A minimum disconnecting set ${\color[rgb]{0,0,0}\{}u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}$ from input $u_{i}=u_{1}$ to all other inputs of parametrized modules is ${\color[rgb]{0,0,0}\{}u_{2}{\color[rgb]{0,0,0}\}}$. In an alternative interpretation it is clear that $u_{2}$ blocks all parallel paths (i.e. parallel to the target module $G_{j1}$) from $u_{i}$ to output $y_{j}$.
The signal $u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$ is composed of $u_{i}=u_{1}$, $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=u_{2}$, while $u_{\!{\scriptscriptstyle\mathcal{T}}}=(u_{3},u_{4}){\color[rgb]{0,0,0}{}^{T}}$. Possible excitation or noise signals $x_{3}$ and $x_{4}$ will serve as $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$. Since $u$ and $e$ are considered to be independent, and $e$ is white noise, the condition (16) can be replaced by a positive definite spectrum of $u_{(1\cup 2)\perp(x_{3},x_{4})}$.
@@
$u_{1}=\sum_{k=1}^{4}F_{1k}(q)x_{k}(t);\ \ \ \ u_{2}=\sum_{k=1}^{4}F_{2k}(q)x_{k}(t).$ (19)
The vector signal $u_{(1\cup 2)\perp(x_{3},x_{4})}$ can then be written as
$u_{(1\cup 2)\perp(x_{3},x_{4})}=\underbrace{\begin{bmatrix}F_{11}(q)&F_{12}(q)\\ F_{21}(q)&F_{22}(q)\end{bmatrix}}_{F(q)}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$
A positive definite spectrum of $u_{(1\cup 2)\perp(x_{3},x_{4})}$ is then obtained for almost all $\omega$ if the matrix $F(z)$ has full row rank over the field of rational functions, and the spectrum of $x_{1}\ \ x_{2}]^{T}$ is positive definite VandenHof&Ramaswamy_CDC:20 .
@@
Moreover the signals $x_{3}$ and $x_{4}$ cannot be used to excite $(u_{1},u_{2})$ since they pass through $u_{\!{\scriptscriptstyle\mathcal{T}}}$ before reaching $u_{2}(=u_{\!{\scriptscriptstyle\mathcal{D}}^{c}})$, and therefore they belong to $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ . To understand why $x_{3}$ and $x_{4}$ cannot be used for excitation of $(u_{1},u_{2})$ we consider the prediction error
$\varepsilon(t,\theta)=H(q,\theta)^{-1}\left[\sum_{k=1}^{4}\Delta G_{jk}(q,\theta)u_{k}(t)+H^{0}(q)e(t)\right].$
Writing $u_{k}$ as a function of the variables $x$, according to the situation of Figure 2, this can be rewritten as
$\varepsilon(t,\theta)=H(q,\theta)^{-1}\left[\sum_{k=1}^{4}T_{k}(q,\theta)x_{k}(t)+H^{0}(q)e(t)\right]$
@@
$T_{1}(q,\theta)$ $=$ $\Delta G_{j1}+\Delta G_{j2}\frac{{\color[rgb]{0,0,0}P}_{21}}{1\!-\!{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}\!+\!\Delta G_{j3}\frac{{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{21}}{1\!-\!{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}$
$T_{2}(q,\theta)$ $=$ $\Delta G_{j2}\frac{1}{1-{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}+\Delta G_{j3}\frac{{\color[rgb]{0,0,0}P}_{32}}{1-{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}$
$T_{3}(q,\theta)$ $=$ $\Delta G_{j2}\frac{{\color[rgb]{0,0,0}P}_{23}}{1-{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}+\Delta G_{j3}\frac{1}{1-{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}$
$T_{4}(q,\theta)$ $=$ $\Delta G_{j2}\frac{{\color[rgb]{0,0,0}P}_{23}{\color[rgb]{0,0,0}P}_{34}}{1\!-\!{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}\!+\!\Delta G_{j3}\frac{{\color[rgb]{0,0,0}P}_{34}}{1\!-\!{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}\!+\!\Delta G_{j4}$
where $\Delta G_{jk}:=G_{jk}^{0}(q)-G_{jk}(q,\theta)$, for $k=1,\cdots 4$. The minimum of the cost function $\bar{\mathbb{E}}\varepsilon^{T}(t,\theta)\varepsilon(t,\theta)$ is achieved if the $T_{k}$-dependent terms are zero. Presence of $x_{k}$ implies that $T_{k}(q,\theta)=0$ in the minimum of the cost function. If $x_{1}$ and $x_{2}$ are present and persistently exciting, then $T_{1}$ and $T_{2}$ are forced to be zero, which implies that $\Delta G_{j1}=0$, and thus a consistent estimate of $G^{0}_{j_{1}}$ is guaranteed. This follows from the observation that $T_{1}(q,\theta)=\Delta G_{j1}+{\color[rgb]{0,0,0}P}_{21}T_{2}(q,\theta)$. If we would replace $x_{2}$ by presence of $x_{3}$, a similar result does not exist. Introducing $x_{4}$ shows that $T_{4}(q,\theta)$ gets involved in the expression and a new unknown term $\Delta G_{j4}$ is introduced also. This explains that forcing $T_{1}$ and $T_{4}$ to be zero is not sufficient for arriving at $\Delta G_{j1}=0$, or in other words: exciting $x_{1}$ and $x_{4}$ does not lead to a consistent estimate of the target module.
In the case that module $G_{j2}(q)$ is not parametrized, but known a priori, there are two options for choosing minimum disconnecting sets:
  • When selecting $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=u_{2}$, it follows that $u_{\!{\scriptscriptstyle\mathcal{T}}}=(u_{3},u_{4}){\color[rgb]{0,0,0}{}^{T}}$, and as a result $x_{3}$ and $x_{4}$ cannot be used for excitation of $(u_{1},u_{2})$, and therefore we need excitations $x_{1}$ and $x_{2}$.
@@
Note that in case multiple disconnecting sets are possible, each one of them provides a set of sufficient conditions for data-informativity. Not satisfying the conditons for one particular choice of disconnecting set, does not necessarily lead to the conclusion that the data-informativity conditions are not satisfied.
In the case that $G_{j4}$ is not parametrized and known a priori, the minimum disconnecting set from $u_{1}$ to ${\color[rgb]{0,0,0}\{}u_{2},u_{3}{\color[rgb]{0,0,0}\}}$ is given by $u_{2}$. Then $u_{\!{\scriptscriptstyle\mathcal{T}}}=u_{3}$, while the external signals $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ that affect $u_{\!{\scriptscriptstyle\mathcal{T}}}$ are given by $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}=(x_{3},x_{4}){\color[rgb]{0,0,0}{}^{T}}$. This implies that (only) signals $(x_{1},x_{2})$ are available for realizing a positive definite spectrum of $(u_{1},u_{2})_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$. Verification of this through the expressions of $T_{k}(q,\theta)$, shows that with $\Delta G_{j4}=0$, the linear combination of the two terms in $T_{3}$ and $T_{4}$ are different from the linear combinations of $\Delta G_{j2}$ and $\Delta G_{j3}$ in the expressions for $T_{1}$ and $T_{2}$. As a result the original result with all modules parametrized follows as the solution.